OpenStudy (anonymous):

the Standard Form of the equation y = −1/3x - 4.

OpenStudy (anonymous):

help!! I think its -1/3+y=-4

OpenStudy (anonymous):

somebody anybody everybody help

OpenStudy (anonymous):

please help!!

OpenStudy (anonymous):

uri! please :(

OpenStudy (anonymous):

ITS: 1/3x + y = -4

OpenStudy (anonymous):

how come its not negative tho

OpenStudy (anonymous):

your right sorry! my answer isnt one of the choices

OpenStudy (anonymous):

To put something in standard form you use this formula: x + y = c So , what do you have to do ..... well , you have to get c alone so you ADD -1/3 on both sides . thats how its not a negative .

OpenStudy (anonymous):

okie dokie ! thanks!

OpenStudy (anonymous):

no probss :)